# Deblurring 3D Gaussian Splatting

### v1

@@ 1 Introduction
Figure 1: Performance comparison to state-of-the-art deblurring NeRFs. Ours achieved a fast rendering speed (200 fps vs. 1 fps) while maintaining competitive rendered image quality (x-axis is represented in log scale).
@@
Recently, 3D Gaussian Splatting (3D-GS) [15] has gained significant attention, demonstrating a capability to produce high-quality images at a remarkably fast rendering speed. It combines a large number of colored 3D Gaussians to represent 3D scenes with a differentiable splatting-based rasterization, substituting NeRF’s time-demanding volumetric rendering. The rendering process projects 3D Gaussian points onto the 2D image planes, and the positions, sizes, rotations, colors, and opacities of these Gaussians are adjusted via gradient-based optimization to better capture the 3D scenes. 3D-GS leverages rasterization, which can be significantly more efficient than volume rendering techniques on modern graphics hardware, thereby enabling rapid real-time rendering.
@@
Expanding on the impressive capabilities of 3D-GS, we aim to further improve its robustness and versatility for more realistic settings, especially those involving blurring effects. Several approaches have attempted to handle the blurring issues in the recent NeRF literature [22, 20, 6, 41, 43]. The pioneering work is Deblur-NeRF [22], which renders sharp images from the neural radiance fields and uses an extra multi-layer perceptron (MLP) to produce the blur kernels. DP-NeRF [20] constrains neural radiance fields with two physical priors derived from the actual blurring process to reconstruct clean images. PDRF [28] uses a two-stage deblurring scheme and a voxel representation to further improve deblurring and training time. All works mentioned above have been developed under the assumption of volumetric rendering, which is not straightforwardly applicable to rasterization-based 3D-GS. Another line of works [41, 6], though not dependent on volume rendering, only address a specific type of blur, i.e., camera motion blur, and are not valid for mitigating the defocus blur.
@@
In this work, we propose Deblurring 3D-GS, the first defocus deblurring algorithm for 3D-GS, which is well aligned with rasterization and thus enables real-time rendering. To do so, we modify the covariance matrices of 3D Gaussians to model the blurriness. Specifically, we employ a small MLP, which manipulates the covariance of each 3D Gaussian to model the scene blurriness. As blurriness is a phenomenon that is based on the intermingling of the neighboring pixels, our Deblurring 3D-GS simulates such an intermixing during the training time. To this end, we designed a framework that utilizes an MLP to learn the variations in different attributes of 3D Gaussians. These small variations are multiplied to the original values of the attributes, which in turn determine the updated shape of the resulting Gaussians. During the inference time, we render the scene using only the original components of 3D-GS without any additional offsets from MLP; thereby, 3D-GS can render sharp images because each pixel is free from the intermingling of nearby pixels. Further, since the MLP is not activated during the inference time, it can still enjoy real-time rendering similar to the 3D-GS while it can reconstruct fine and sharp details from the blurry images.
Figure 2: Our method’s overall workflow. $\gamma(\cdot)$ denotes positional encoding, $\odot$ denotes hadamard product, and $x$, $r$, $s$ stand for position, quaternion, and scaling of 3D Gaussian respectively. Given a set of 3D Gaussians $G$, we extract $x$, $r$, $s$ from each 3D Gaussian. Viewing direction $v$ from the dataset and $x$ are positionally encoded and then fed to $F_{\theta}$, a small MLP parameterized with $\theta$, which yields offsets $\delta r$ and $\delta s$ for each Gaussian. These $\delta r$ and $\delta s$ are element-wisely multiplied to their respective Gaussians’s $r$ and $s$ respectively, and these computed attributes are used to construct transformed 3D Gaussians $G^{\prime}(x,r\cdot\delta r,s\cdot\delta s)$. These transformed Gaussians $G^{\prime}$ are differentiably rasterized to render blurred image during training time. However, for inference stage, $G$ is directly fed to differentiable rasterizer, without involving any MLP, to render sharp image.
@@
3D-GS [15] models a 3D scene from a sparse point cloud, which is usually obtained from the structure-from-motion (SfM) [35]. SfM extracts features from multi-view images and relates them via 3D points in the scene. If the given images are blurry, SfM fails heavily in identifying the valid features, and ends up extracting a very small number of points. Even worse, if the scene has a larger depth of field, SfM hardly extracts any points which lie on the far end of the scene. Due to this excessive sparsity in the point cloud constructed from set of blurry images, existing methods, including 3D-GS [15], that rely on point clouds fail to reconstruct the scene with fine details. To compensate for this excessive sparsity, we propose to add extra points with valid color features to the point cloud using N-nearest-neighbor interpolation [29]. In addition, we prune Gaussians based on their position to keep more Gaussians on the far plane.
@@
A variety of experiments have been conducted on the benchmark, and the results have revealed the effectiveness of our approach for deblurring. Tested under different evaluation matrices, our method achieves state-of-the-art rendering quality or performs on par with the currently leading models while achieving significantly faster rendering speed ($>$ 200 FPS)
@@
  • We propose the first real-time rendering-enabled defocus deblurring framework using 3D-GS.
  • We propose a novel technique that manipulates the covariance matrix of each 3D Gaussian differently to model spatially changing blur using a small MLP.
@@
  • We achieve FPS $>$ 200 while accomplishing superior rendering quality or performing on par with the existing cutting-edge models under different metrics.
@@ 2.1 Image Deblurring
It is common to observe that when we casually take pictures with optical imaging systems, some parts of scene appear blurred in the images. This blurriness is caused by a variety of factors, including object motion, camera shake, and lens defocusing [1, 34]. When the image plane is separated from the ideal reference plane during the imaging process, out-of-focus or defocus blur happens. The degradation induced by defocus blur of an image is generally expressed as follows:
@@
The first canonical approach to deconvolve an image is [33]. It applies iterative minimization to an energy function to obtain a maximum likelihood approximation of the original image, using a known PSF (the blur kernel). Typically, the blurring kernel $K$ and the sharp image $I$ are unknown. The technique termed as image blind deblurring [42] involves recovering the latent sharp image given just one blurry image $B$. In this procedure, the deblurred image is obtained by first estimating a blurring kernel. Image blind deblurring is a long-standing and ill-posed problem in the field of image and vision. Numerous approaches to address the blurring issue have been suggested in the literature. Among them, traditional methods often construct deblurring as an optimization problem and rely on natural image priors [21, 26, 44, 47]. Conversely, the majority of deep learning-based techniques use convolutional neural networks (CNN) to map the blurry image with the latent sharp image directly [25, 46, 31]. While series of studies have been actively conducted for image deblurring, they are mainly designed for deblurring 2D images and are not easily applicable to 3D scenes deblurring due to the lack of 3D view consistency.
@@ 2.2 Neural Radiance Fields
Neural Radiance Fields (NeRF) are a potent method that has recently gained popularity for creating high-fidelity 3D scenes from 2D images. Realistic rendering from novel viewpoints is made possible by NeRF, which uses deep neural networks to encode volumetric scene features. To estimate density $\sigma\in[0,\infty)$ and color value $c\in[0,1]^{3}$ of a given point, a radiance field is a continuous function $f$ that maps a 3D location $x\in\mathbb{R}^{3}$ and a viewing direction $d\in\mathbb{S}^{2}$. This function has been parameterized by a multi-layer perceptron (MLP) [23], where the weights of MLP are optimized to reconstruct a series of input photos of a particular scene: $(c,\sigma)=f_{\theta}:\big{(}\gamma(x),\gamma(d)\big{)}$ Here, $\theta$ indicates the network weights, and $\gamma$ is the specified positional encoding applied to $x$ and $d$ [39]. To generate the images at novel views, volume rendering [7] is used, taking into account the volume density and color of points.
@@ 2.2.1 Fast Inference NeRF
Numerous follow-up studies have been carried out to enhance NeRF’s rendering time to achieve real-time rendering. Many methods, such as grid-based approaches [40, 4, 38, 3, 8, 32], or those relying on hash [24, 2] adopt additional data structures to effectively reduce the size and number of layers of MLP and successfully improve the inference speed. However, they still fail to reach real-time view synthesis. Another line of works [30, 13, 45] proposes to bake the trained parameters into the faster representation and attain real-time rendering. While these methods rely on volumetric rendering, recently, 3D-GS [15] successfully renders photo-realistic images at novel views with noticeable rendering speed using a differentiable rasterizer and 3D Gaussians. Although several approaches have attempted to render tens or hundreds of images in a second, deblurring the blurry scene in real-time is not addressed, while blurriness commonly hinders clean image acquisition in the wild.
@@ 2.2.2 Deblurring NeRF
Several strategies have been proposed to train NeRF to render clean and sharp images from blurry input images. While DoF-NeRF [43] attempts to deblur the blurry scene, both all-in-focus and blurry images are required to train the model. Deblur-NeRF [22] firstly suggests deblurring NeRF without any all-in-focus images during training. It employs an additional small MLP, which predicts per-pixel blur kernel. However, it has to render the neighboring pixels of the target pixel in advance to model the interference of the neighboring information, so it takes additional computation at the time of training. Though the inference stage does not involve the blur kernel estimation, it is no different from the training with regard to rendering time as it is based on volumetric rendering which takes several seconds to render a single image. DP-NeRF [19] improved Deblur-NeRF in terms of image quality, and PDRF [28] partially replaces MLP to the grid and ameliorates rendering speed, still they depend on volumetric rendering to produce images and are not free from the rendering cost. Other works [6, 41] are aimed at addressing another type of blur, camera motion blur, rather than solving the long rendering time. While these deblurring NeRFs successfully produce clean images from the blurry input images, there is room for improvement in terms of rendering time. Thus, we propose a novel deblurring framework, deblurring 3D Gaussian Splatting, which enables real-time sharp image rendering using a differentiable rasterizer and 3D Gaussians.
@@ 3 Deblurring 3D Gaussian Splatting
Based on the 3D-GS [15], we generate 3D Gaussians, and each Gaussian is uniquely characterized by a set of the parameters, including 3D position $x$, opacity $\sigma$, and covariance matrix derived from quaternion $r$ scaling $s$. Each 3D Gaussian also contains spherical harmonics (SH) to represent view-dependent appearance. The input for the proposed method consists of camera poses and point clouds, which can be obtained through the structure from motion (SfM) [35], and a collection of images (possibly blurred). To deblur 3D Gaussians, we employ an MLP that takes the positions of the 3D Gaussians along with the $r$ and $s$ as inputs and outputs $\delta r$, and $\delta s$ which are the small scaling factors multiplied to $r$ and $s$, respectively. With new quaternion and scale, $r\cdot\delta r$ and $s\cdot\delta s$, the updated 3D Gaussians are subsequently fed to the tile-based rasterizer to produce the blurry images. The overview of our method is shown in Fig. 2.
@@ 3.2.1 Motivation
It is discussed in eq. 1 that the pixels in images get blurred due to defocusing, and this phenomenon is usually modeled through a convolution operation. According to the thin lens law [12], the scene points that lie at the focal distance of the camera make a sharp image at the imaging plane. On the other hand, any scene points that are not at the focal distance will make a blob instead of a sharp point on the imaging plane, and it produces a blurred image. If the separation from the focal distance of a scene point is large, it produces a blob of a large area, which corresponds to severe blur. This blurring effect is usually modeled by a 2D Gaussian function known as the point spread function (PSF). Correspondingly, an image captured by a camera is the result of the convolution of the actual image and the PSF. Through this convolution, which is the weighted summation of neighboring pixels, few pixels can affect the central pixel heavily depending on the weight. In other words, in the defocus imaging process, a pixel affects the intensity of neighboring pixels. This theoretical base provides us the motivation to build our deblurring 3D Gaussians framework. We assume that big sized 3D Gaussians cause the blur, while relatively smaller 3D Gaussians correspond to the sharp image. This is because those with greater dispersion are affected by more neighboring information as they are responsible for wider regions in image space, so they can represent the interference of the neighboring pixels. Whereas the fine details in the 3D scene can be better modeled through the smaller 3D Gaussians.
Figure 3: Comparison to normal Gaussian blur kernel. Top row: It shows the proposed method. $g$ is the Gaussian before the transformation, and $g^{\prime}$ is the Gaussian after the transformation. Since the different transformations can be applied to different Gaussian, ours can selectively blur the images depending on the scene; it can only blur the front parts of the scene. Bottom row: It shows a normal Gaussian blur kernel, which is not capable of handling different parts of the image differently, thereby uniformly blurring the entire image.
3.2.2 Modelling
Following the above mentioned motivation, we learn to deblur by transforming the geometry of the 3D Gaussians. The geometry of the 3D Gaussians is expressed through the covariance matrix, which can be decomposed into the rotation and scaling factors as mentioned in Eq. 3. Therefore, our target is to change the rotation and scaling factors of 3D Gaussians in such a way that we can model the blurring phenomenon. To do so, we have employed an MLP that takes the position $x$, rotation $r$, scale $s$, and viewing direction $v$ of 3D Gaussians as inputs, and outputs $(\delta r,\delta s)$, as given by:
$(\delta r,\delta s)=\mathcal{F_{\theta}}\Big{(}\gamma(x),r,s,\gamma(v)\Big{)},\vspace{0.2cm}$ (6)
@@ 3.2.2 Modelling
The minima of these scaling factors $(\delta r,\delta s)$ are clipped to 1 and element-wisely multiplied to $r$ and $s$, respectively, to obtain the transformed attributes $r^{\prime}=r\cdot\delta r$ and $s^{\prime}=s\cdot\delta s$. With these transformed attributes, we can construct the transformed 3D Gaussians $G(x,r^{\prime},s^{\prime})$, which is optimized during training to model the scene blurriness. As $s^{\prime}$ is greater than or equal to $s$, each 3D Gaussian of $G(x,r^{\prime},s^{\prime})$ has greater statistical dispersion than the original 3D Gaussian $G(x,r,s)$. With the expanded dispersion of 3D Gaussian, it can represent the interference of the neighboring information which is a root cause of defocus blur. In addition, $G(x,r^{\prime},s^{\prime})$ can model the blurry scene more flexibly as per-Gaussian $\delta r$ and $\delta s$ are estimated. Defocus blur is spatially-varying, which implies different regions has different levels of blurriness. The scaling factors for 3D Gaussians that are responsible for a region with harsh defocus blur where various neighboring information in wide range is involved in, becomes bigger to better model high degree of blurriness. Those for 3D Gaussians on the sharp area, on the other hand, are closer to 1 so that they have smaller dispersion and do not represent the influence of the nearby information.
@@
At the time of inference, we use $G(x,r,s)$ to render the sharp images. As mentioned earlier, we assume that multiplying two different scaling factors to transform the geometry of 3D Gaussians can work as blur kernel and convolution in eq. 1. Thus, $G(x,r,s)$ can produce the images with clean and fine details. It is worth noting that since any additional scaling factors are not used to render the images at testing time, $F_{\theta}$ is not activated so all steps required for the inference of Deblurring 3D-GS are identical to 3D-GS, which in turn enables real-time sharp image rendering. In terms of training, only forwarding $F_{\theta}$ and simple element-wise multiplication are extra cost.
3.2.3 Selective blurring
The proposed method can handle the training images arbitrarily blurred in various parts of the scene. Since we predict $(\delta r,\delta s)$ for each Gaussian, we can selectively enlarge the covariances of Gaussians where the parts in the training images are blurred. Such transformed per-Gaussian covariance is projected to 2D space and it can act as pixel-wise blur kernel at the image space. It is noteworthy that applying different shapes of blur kernels to different pixels plays a pivotal role in modeling scene blurriness since blurriness spatially varies. This flexibility enables us to effectively implement deblurring capability in 3D-GS. On the other hand, a naive approach to blurring the rendered image is simply to apply a Gaussian kernel. As shown in Fig. 3, this approach will blur the entire image, not blur pixel-wisely, resulting in blurring parts that should not be blurred for training the model. Even if learnable Gaussian kernel is applied, optimizing the mean and variance of the Gaussian kernel, a single type of blur kernel is limited in its expressivity to model the complicatedly blurred scene and is optimized to model the average blurriness of the scene from averaging loss function which fails to model blurriness morphing in each pixel. Not surprisingly, the Gaussian blur is a special case of the proposed method. If we predict one $(\delta s)$ for all 3D Gaussians, then it will similarly blur the whole image.

### v2

@@ Abstract
Keywords:
Neural Radiance Fields Deblurring Real-time rendering
@@ 1 Introduction
Figure 1: Performance comparison to state-of-the-art deblurring NeRFs. Ours achieved a fast rendering speed ($>$ 800 FPS vs. 1 FPS) while maintaining competitive rendered image quality (the x-axis is represented in log scale).
@@
Recently, 3D Gaussian Splatting (3D-GS) [14] has gained significant attention, demonstrating a capability to produce high-quality images at a remarkably fast rendering speed. Substituting NeRF’s time-demanding volumetric rendering, it combines a large number of colored 3D Gaussians to represent 3D scenes with a differentiable splatting-based rasterization, which can be significantly more efficient than volume rendering techniques on modern graphics hardware, thereby enabling rapid real-time rendering.
@@
Expanding on the impressive capabilities of 3D-GS, we aim to further improve its robustness and versatility for more realistic settings, especially those involving blurring effects. Several approaches have attempted to handle the blurring issues in the recent NeRF literature [22, 20, 5, 39, 40]. The pioneering work is Deblur-NeRF [22], which renders sharp images from images with defocus blur or camera motion blur using an extra multi-layer perceptron (MLP) to produce the blur kernels. DP-NeRF [20] constrains neural radiance fields with two physical priors derived from the actual blurring process to reconstruct clean images. PDRF [28] uses a two-stage deblurring scheme and a voxel representation to further improve deblurring and training time. All works mentioned above have been developed under the assumption of volumetric rendering, which is not straightforwardly applicable to rasterization-based 3D-GS. Another line of works [39, 5, 18] though not dependent on volume rendering, only address a single specific type of blur, i.e., either camera motion blur or defocus blur, and are not valid for mitigating the both types of blur.
@@
In this work, we propose Deblurring 3D-GS, the first deblurring algorithm for 3D-GS, which is well aligned with rasterization and thus enables real-time rendering. To do so, we modify the covariance matrices of 3D Gaussians to model the blurriness. Specifically, we employ a small MLP, which manipulates the covariance mean of each 3D Gaussian to model the scene blurriness. As blurriness is a phenomenon that is based on the intermingling of the neighboring pixels, our Deblurring 3D-GS simulates such an intermixing during the training time. To this end, we designed a framework that utilizes an MLP to learn the variations in different attributes of 3D Gaussians. These small variations are multiplied or added to the original values of the attributes, which in turn determine the updated shape of the resulting Gaussians. During the inference time, we render the scene using only the original components of 3D-GS without any additional outputs from the MLP; thereby, 3D-GS can render sharp images because each pixel is free from the intermingling of nearby pixels. Further, since the MLP is not activated during the inference time, it can still enjoy real-time rendering similar to the 3D-GS while it can reconstruct fine and sharp details from the blurry images.
Figure 2: Our method’s overall workflow. $\gamma(\cdot)$ denotes positional encoding, $\odot$ and $\oplus$ denotes hadamard product and averaging operation for each, and $x$, $r$, $s$ stand for position, quaternion, and scaling of 3D Gaussian respectively. $\otimes$ is an operator that implements $\delta r\odot r$, $\delta s\odot s$, and $\delta x+x$. Dotted arrows and dashed arrows describe the pipeline for modeling camera motion blur and modeling defocus blur, respectively at training time. Solid arrows show the process of rendering sharp images at the inference time. More details are explained at  Sec. 3.2.
@@
3D-GS [14] models a 3D scene from a sparse point cloud, which is usually obtained from the structure-from-motion (SfM) [34]. SfM extracts features from multi-view images and relates them via 3D points in the scene. If the given images are blurry, SfM fails heavily in identifying the valid features, and ends up extracting a very small number of points. Even worse, if the scene has a larger depth of field, SfM hardly extracts any points which lie on the far end of the scene. Due to this excessive sparsity in the point cloud constructed from set of blurry images, existing methods, including 3D-GS [14], that rely on point clouds fail to reconstruct the scene with fine details. To compensate for this excessive sparsity, we propose to add extra points with valid color features to the point cloud using K-nearest-neighbor interpolation [29]. In addition, we prune Gaussians based on their position to keep more Gaussians on the far plane.
@@
A variety of experiments have been conducted on the benchmark, and the results have revealed the effectiveness of our approach for deblurring. Tested under different evaluation matrices, our method achieves state-of-the-art rendering quality or performs on par with the currently leading models while achieving significantly faster rendering speed ($>$ 800 FPS)
@@
  • We propose the first real-time rendering deblurring framework using 3D-GS.
  • We propose a novel technique that manipulates the covariance matrix and mean of each 3D Gaussian differently to model spatially changing blur using a small MLP.
@@
  • We achieve FPS $>$ 800 while accomplishing superior rendering quality or performing on par with the existing cutting-edge models under different metrics.
@@ 2.1 Image Deblurring
It is common to observe that when we casually take pictures with optical imaging systems, some parts of the scene appear blurred in the images. This blurriness is caused by a variety of factors, including object motion, camera shake, and lens defocusing [1, 33]. The degradation induced by the blur of an image is generally expressed as follows:
@@
Traditional methods often construct deblurring as an optimization problem and rely on natural image priors [21, 27, 41, 45]. Conversely, the majority of deep learning-based techniques use convolutional neural networks (CNN) to map the blurry image with the latent sharp image directly [26, 44, 31]. While a series of studies have been actively conducted for image deblurring, they are mainly designed for deblurring 2D images and are not easily applicable to 3D scenes deblurring due to the lack of 3D view consistency.
@@ 2.2 Neural Radiance Fields
Neural Radiance Fields (NeRF) is a potent method that has gained popularity for creating high-fidelity 3D scenes from 2D images, employing deep neural networks to encode volumetric scene features. To estimate density $\sigma\in[0,\infty)$ and color value $c\in[0,1]^{3}$ of a given point, a radiance field is a continuous function $f$ that maps a 3D location $x\in\mathbb{R}^{3}$ and a viewing direction $d\in\mathbb{S}^{2}$. This function has been parameterized by a multi-layer perceptron (MLP) [23], where the weights of MLP are optimized to reconstruct a series of input photos of a particular scene: $(c,\sigma)=f_{\theta}:\big{(}\gamma(x),\gamma(d)\big{)}$. Here, $\theta$ indicates the network weights, and $\gamma$ is the specified positional encoding applied to $x$ and $d$ [37]. To generate the images at novel views, volume rendering [6] is used, taking into account the volume density and color of points.
@@ 2.2.1 Fast Inference NeRF
Numerous follow-up studies have been carried out to enhance NeRF’s rendering time to achieve real-time rendering. Many methods, such as grid-based approaches [38, 4, 36, 3, 8, 32, 25], or those relying on hash [24, 2] adopt additional data structures to effectively reduce the size and number of layers of MLP and successfully improve the inference speed. However, they still fail to reach real-time view synthesis. Another line of works [30, 13, 42] proposes to bake the trained parameters into the faster representation and attain real-time rendering. While these methods rely on volumetric rendering, recently, 3D-GS [14] successfully renders photo-realistic images at novel views with noticeable rendering speed using a differentiable rasterizer and 3D Gaussians. Although several approaches have attempted to render tens or hundreds of images in a second, deblurring the blurry scene in real-time is not addressed, while blurriness commonly hinders clean image acquisition in the wild.
@@ 2.2.2 Deblurring NeRF
Several strategies have been proposed to train NeRF to render clean and sharp images from blurry input images. While DoF-NeRF [40] attempts to deblur the blurry scene, both all-in-focus and blurry images are required to train the model. Deblur-NeRF [22] firstly suggests deblurring NeRF without any all-in-focus images during training. It employs an additional small MLP, which predicts per-pixel blur kernel to model defocus and camera motion blur. Though the inference stage does not involve the blur kernel estimation, it is no different from the training with regard to rendering time as it is based on volumetric rendering which takes several seconds to render a single image. DP-NeRF [19] and PDRF [28] further improved Deblur-NeRF, still they depend on volumetric rendering and are not free from the rendering cost. Other approaches [5, 39, 18] are bounded to addressing only one type of blur, either camera motion blur or defocus blur, and not aimed at solving the long rendering time. While these deblurring NeRFs successfully produce clean images from the blurry input images, there is room for improvement in terms of rendering time. Thus, we propose a novel deblurring framework, Deblurring 3D Gaussian Splatting, which enables real-time sharp image rendering using a differentiable rasterizer and 3D Gaussians.
@@ 3 Deblurring 3D Gaussian Splatting
Based on the 3D-GS [14], we generate 3D Gaussians, and each Gaussian is uniquely characterized by a set of the parameters, including 3D position $x$, opacity $\sigma$, and covariance matrix derived from quaternion $r$ scaling $s$. Every 3D Gaussian also contains spherical harmonics (SH) to represent view-dependent appearance. The input for the proposed method consists of camera poses and point clouds, which can be obtained through the structure from motion (SfM) [34], and a collection of images (possibly blurred). We employ an MLP that takes $x_{j}$, $r_{j}$, and $s_{j}$ which are 3D position, quaternion, and scaling of $j$-th Gaussian, respectively as inputs to deblur a scene. In case of modeling defocus blur, the MLP yields $\delta r_{j}$ and $\delta s_{j}$ which are the small scaling factors multiplied to $r$ and $s$, respectively. With new quaternion and scale, $r_{j}\cdot\delta r_{j}$ and $s_{j}\cdot\delta s_{j}$, the updated 3D Gaussians are subsequently fed to the tile-based rasterizer to rasterize the defocus blurred images. To address camera motion blur, the MLP outputs $\{(\delta x_{j}^{(i)},\delta r_{j}^{(i)},\delta s_{j}^{(i)})\}_{i=1}^{M}$ where $M$ is the number of the auxiliary sets of 3D Gaussians representing the moments of camera movement and $\delta x_{j}^{(i)},\delta r_{j}^{(i)},\delta s_{j}^{(i)}$ are $i$-th predicted position offset, scaling factor for scale, and scaling factor for quaternion of $j$-th 3D Gaussian, respectively. Rasterizer produces $M$ images from $M$ different sets of 3D Gaussians and we average them to obtain camera motion blurred image. The overview of our method is shown in Fig. 2.
@@ 3.2.1 Motivation
It is discussed in Eq. 1 that the pixels in images get blurred due to defocusing and camera motion, and this phenomenon is usually modeled through a convolution operation. Correspondingly, an image captured by a camera is the result of the convolution of the actual image and the PSF. Through convolution, which is the weighted summation of neighboring pixels, some pixels can affect the central pixel heavily depending on the weight. In other words, in the blurry imaging process, a pixel affects the intensity of neighboring pixels. This theoretical base motivates us to build our Deblurring 3D Gaussians framework.
When handling defocus blur, we assume that big-sized 3D Gaussians cause the blur, while smaller 3D Gaussians correspond to the sharp image. This is because those with greater dispersion are affected by more neighboring information as they are responsible for wider regions in image space, so they can represent the interference of the neighboring pixels. Whereas the fine details in the 3D scene can be better modeled through the smaller 3D Gaussians. In the case of camera motion blur, we implicitly model the camera movement during the camera exposure time. In detail, we generate multiple auxiliary sets of 3D Gaussians which represent the discrete moment of the movement, by shifting the positions of the existing set of 3D Gaussians, and simulate camera motion blur. More details are described in the supplementary material.
3.2.2 Defocus blur modeling
Following the aforementioned motivation, we learn to deblur by transforming the geometry of the 3D Gaussians. The geometry of the 3D Gaussians is expressed through the covariance matrix, which can be decomposed into the rotation and scaling factors as mentioned in Eq. 3. Therefore, our target is to change the rotation and scaling factors of 3D Gaussians in such a way that we can model the blurring phenomenon. To do so, we have employed an MLP that takes the position $x_{j}$, rotation $r_{j}$, scale $s_{j}$ of $j$-th 3D Gaussian, and viewing direction $v$ as inputs, and outputs $(\delta r_{j},\delta s_{j})$, as given by:
$(\delta r_{j},\delta s_{j})=\mathcal{F_{\theta}}\Big{(}\gamma(x_{j}),r_{j},s_{j},\gamma(v)\Big{)},\vspace{0.2cm}$ (6)
@@ 3.2.2 Defocus blur modeling
Each scaling factor $(\delta r_{j},\delta s_{j})$ is scaled by $\lambda_{s}$, shifted by $(1-\lambda_{s})$ for optimization stability. Then the minima of them are clipped to 1.0 and element-wisely multiplied to $r_{j}$ and $s_{j}$, respectively, to obtain the transformed attributes as following:
$\hat{r}_{j}=r_{j}\cdot\text{min}(1.0,~{}\lambda_{s}\delta r_{j}+(1-\lambda_{s})),$ (8)
$\hat{s}_{j}=s_{j}\cdot\text{min}(1.0,~{}\lambda_{s}\delta s_{j}+(1-\lambda_{s})),$ (9)
@@
where function $\text{min}(\cdot,\cdot)$ returns smaller value. With these transformed attributes, we can construct the transformed 3D Gaussians $G(x_{j},\hat{r}_{j},\hat{s}_{j})$, which is optimized during training to model the scene blurriness. As $\hat{s}_{j}$ is greater than or equal to $s$, each 3D Gaussian of $G(x_{j},\hat{r}_{j},\hat{s}_{j})$ has greater statistical dispersion than the original 3D Gaussian $G(x_{j},r_{j},s_{j})$. With the expanded dispersion of 3D Gaussian, it can represent the interference of the neighboring information which is a root cause of defocus blur. In addition, $G(x_{j},\hat{r}_{j},\hat{s}_{j})$ can model the blurry scene more flexibly as per-Gaussian $\delta r$ and $\delta s$ are estimated. Defocus blur is spatially varying, which implies different regions have different levels of blurriness. The scaling factors for 3D Gaussians that are responsible for a region with harsh defocus blur where various neighboring information in wide range is involved in, become bigger to better model a high degree of blurriness. Meanwhile, those for 3D Gaussians on the sharp area are closer to 1.0 so that they have smaller dispersion and do not represent the influence of the nearby information. Therefore, we can model defocus blur and rasterize defocus blurred image with $G(x_{j},\hat{r}_{j},\hat{s}_{j})$.
@@
At the time of inference, we use $G(x_{j},r_{j},s_{j})$ to render the sharp images. As mentioned earlier, we assume that multiplying two different scaling factors to transform the geometry of 3D Gaussians can work as blur kernel and convolution in Eq. 1. Thus, $G(x_{j},r_{j},s_{j})$ can produce the images with clean and fine details. It is worth noting that since any additional scaling factors are not used to render the images at testing time, $F_{\theta}$ is not activated, so all steps required for the inference of Deblurring 3D-GS are identical to 3D-GS, which in turn enables real-time sharp image rendering.
3.2.3 Camera motion blur modeling
We model camera motion blur with additional sets of 3D Gaussians. We adjust the geometry of each 3D Gaussian to simulate blur at training time, akin to defocus blur modeling. However, unlike defocus blur, camera motion blur occurs due to the physical movement of a camera. Every moment when the light hits the camera sensor, camera movement during the exposure time makes light intensities from multiple sources intermixed. We model such a phenomenon by adding small offsets to the position of each 3D Gaussian, and produce additional sets of 3D Gaussians to implicitly represent camera shake, and average clean images from different moments to simulate camera motion blur. Specifically, we first slightly change Eq. 6 to compute additional output $\delta x$, the offset for the position $x_{j}$ of a Gaussian, from $F_{\theta}$ as:
$\{(\delta x_{j}^{(i)},\delta r_{j}^{(i)},\delta s_{j}^{(i)})\}_{i=1}^{M}=\mathcal{F_{\theta}}\Big{(}\gamma(x_{j}),r_{j},s_{j},\gamma(v)\Big{)},\vspace{0.2cm}$ (10)
where $M$ is the number of additional sets of 3D Gaussian to model the moments during the camera movement, $\delta x_{j}^{(i)}$ is $i$-th predicted position offset of $j$-th Gaussian. $\delta r_{j}^{(i)}~{}\text{and}~{}\delta s_{j}^{(i)}$ are scaled, shifted, and clipped in the same manner to Eq. 8 and Eq. 9 for each. Then we construct extra $M$ sets of 3D Gaussians $\{\{(\hat{x}_{j}^{(i)},\hat{r}_{j}^{(i)},\hat{s}_{j}^{(i)})\}_{i=1}^{M}\}_{j=1}^{N_{G}}$ by shifting the positions and changing the geometry of the existing set of 3D Gaussians, where $N_{G}$ is the number of the current 3D Gaussians, $\hat{x}_{j}^{(i)}$ stands for the shifted position with scaled $\delta x_{j}^{(i)}$ by $\lambda_{p}$. $\hat{x}_{j}^{(i)}=x_{j}+\lambda_{p}\delta x_{j}^{(i)}$, and $\hat{r}_{j}^{(i)}=r_{j}\cdot\delta r_{j}^{(i)}~{}\text{and}~{}\hat{s}_{j}^{(i)}=s_{j}\cdot\delta s_{j}^{(i)}$ are computed in a similar way to compute them for defocus blur deblurring. Each set corresponds to 3D Gaussians observed from different camera viewpoints and we rasterize $M$ clean images from $M$ different sets of 3D Gaussians and then average them to obtain a single camera motion blurred image $I_{b}$ at the training time as following:
$I_{b}=\frac{1}{M}\sum_{i=1}^{M}I_{i},\quad I_{i}=\texttt{Rasterize}(\{G(\hat{x}_{j}^{(i)},\hat{r}_{j}^{(i)},\hat{s}_{j}^{(i)})\}_{j=1}^{N_{G}}),$ (11)
where $I_{i}$ is a clean image generated by the predicted deltas. Deblurring camera motion blur also does not require any MLP forwarding and rendering multiple images at the inference time as $G(x_{j},r_{j},s_{j})$ learns the latent clean image. Thus, we can still enjoy rendering clean images from camera motion blurred input images in real-time manners, just like defocus blur deblurring.